A ball is raised above the floor such that it has 5.0 J of gravitational potential energy and then allowed to fall. How much kinetic energy will it have just before striking the floors?
1
Expert's answer
2020-12-08T10:52:45-0500

According to the law conservation of energy, the total energy of an isolated system remains constant.

The total energy consists of the sum of potential energy "U" and kinetic energy "K":


"E = U+K = mgh + \dfrac{mv^2}{2} = const"

where "h" is the height of the ball, "v" is its speed and "m" is its mass. Since initially the ball didn't have any speed "v = 0", the total energy is equal to the potential one.


"E = U_0 = 5J"

On the opposite just before striking the floors the ball didn't have any potential energy ("h = 0" ), thus, its kinetic energy at that moment was equal to the total energy "E" (that is constant at any time):


"K = E=5J"

Answer. 5J.